CRY+VIS Interface Meeting
Date: 2022/5/20 13:00-
Review of IY and EY
Date: 2022/5/20 14:00-14:20
Participants: Takahashi, Fabian, Uchiyama, Washimi, Ushiba, Aso, Tanaka, Tomaru
- Disassembled the footholds around the top chamber in EY.
- Review of IY and EY.
- SR2 geophone calibration: I found an error in my script, and now I'm getting more reasonable results. See the klog.
- In SR2, I calculated inter-calibration factors between geophones and LVDTs. See the klog for details.
- Measured the transfer functions of the PR suspensions.
- We installed new cable clamp boards, and clamped the cables such that no connectors were hanging.
- We installed the limit switches for the Traverser in PRM.
- We still need to connect them to the drivers, but we don't have the drivers yet.
- We should check the health of the cables we clamped on the new boards. We don't have a date to do this work yet.
- We should decide a date to test the limit switches, but I guess that depends on when the drivers will become available.
- Installed the repaired high power coil driver to the OMMT1.
- Type-A: Monitoring of the evacuation in IY/EY.
- Type-B/Bp/C: Calibration in SR. FC for BS. Cable check in PR. Installation of OMC.
(Date in Kamioka)
The week from 5/23
- Takahashi: 5/23-26
- Sato: 5/24-27
- Hirata: 5/23-27
- Aso: 5/23-26
The week from 5/30
- Takahashi: 5/30-6/3
- Sato: 5/30-6/3
- Hirata: 5/30-6/3